A certain weak base has a Ky of 8.50 x 10-7. What concentration of this base will produce a pH of 10.05? concentration:
Answer #1

kn- dissociation Constant of weakbase, Given, kb = 8.50xlo? Given, pH of Solution -10.05. pH + pOH = 14 pºn. 14-p14-100 1 =Сон-) = С. Ec. The SENSE NE Сонг2 = укс 1-12* 10% = J 8. soxc0²x (1+12 x16**. 8. S9 x tsºxc c = (1.12*104)? = 8.50X167 = 1.25